mathslover (mathslover):

Indefinite Integral \(\large{ \int \sin (2x) \cos (3x) dx }\)

ganeshie8 (ganeshie8):

2sinAcosB identity

mathslover (mathslover):

\(\large{ \int 2\sin x \cos x \left[4 \cos^3 (x) - 3\cos (x) \right] }\)

ganeshie8 (ganeshie8):

this identity simplifies the work : 2sinAcosB = sin(A+B) + sin(A-B)

mathslover (mathslover):

\(\large{ 2 \int \sin x \cos x (\cos x ) \left[ 4\cos^2 (x) - 3 \right] \\ 2 \int \sin x \cos^2 x \left[ 4(1-\sin^2 x) -3 \right] }\)

mathslover (mathslover):

\(\large{ 2 \int \sin x ( 1- \sin^2 x) \left[ 1 - 4\sin^2 x \right] }\) Am I going on wrong track?

ganeshie8 (ganeshie8):

nope, you're in right track... but thats a lengthy procss

mathslover (mathslover):

Yeah.. that seems to be lengthy. I tried a lot, though, I didn't get a feeling that I am on wrong track, but the answer I got doesn't match the correct answer.

ganeshie8 (ganeshie8):

try working it again using the angle sum formula

mathslover (mathslover):

Using the angle sum formula here? Never did this before, can you give me a brief idea for this?

mathslover (mathslover):

Okay, well, I think I got your point : \(\int \sin (2x) \cos (3x) \) \(\cfrac{1}{2} \int 2\sin (2x) \cos (3x) \)

mathslover (mathslover):

\(\cfrac{1}{2} \int (\sin (5x) - \sin (x) )dx\)

ganeshie8 (ganeshie8):

Yep !

mathslover (mathslover):

Okay, yeah, am getting that @ganeshie8 ...!

mathslover (mathslover):

So, that becomes : \(\cfrac{1}{2} \times \cfrac{-\cos^5 x}{5} - \cfrac{1}{2} \times (-\cos x) +C\)

ganeshie8 (ganeshie8):

Looks good !

ganeshie8 (ganeshie8):

oh wait a sec, bring that exponent back to its place lol

mathslover (mathslover):

Got it... \(\cfrac{-1}{10} (\cos 5 x) + \cfrac{1}{2} \cos x + C\) Oh sorry lol

mathslover (mathslover):

That 5 was so nasty... didn't want to live with x... so, he decided to become an exponent! I was helpless :P Jokes apart, Thanks a lot @ganeshie8 .. You did it again. :)

Miracrown (miracrown):

oh hahaa

OpenStudy (vishweshshrimali5):

""The nasty5"" by @mathslover. First day first show at OS Talkies. Hurry before the film becomes HOUSE FULL lol ;) :P

mathslover (mathslover):

Ah... the ticket costs 1 billion dollars. Who wants to buy it?

OpenStudy (vishweshshrimali5):

For tickets in black contact @vishweshshrimali5. Hurry offer limited.

OpenStudy (vishweshshrimali5):

Well, it seems the people are not that interested in nasty5 we will have to work for the sequel. LOL

mathslover (mathslover):

Flop show :( May be @ganeshie8 and @Miracrown want to buy the ticket? (0.1 % discount for them :P )

ganeshie8 (ganeshie8):

lol the talk in the town is that they're giving the tickets for free :P

Miracrown (miracrown):

nty, films don't interest me.

OpenStudy (vishweshshrimali5):

Well we also have the comic version of NASTY 5. Signed by superman and batman. @Miracrown Interested ? @mathslover is bearing the whole cost .

OpenStudy (vishweshshrimali5):

@ganeshie8 I started the rumour.

OpenStudy (vishweshshrimali5):

Ohh no, I spoiled the serious environment. Bye guys ;) And yes great work @ganeshie8 and @mathslover

mathslover (mathslover):

Someone just stole my piggy bank... It contained 1 billion dollars. Anyone interested to donate ? =P Well, being an ambi, I am bound to abide with the rules... :( This has become off-topic, so we will continue this discussion when we get the answer in the next post. Thanks @ganeshie8 and its okay @vishweshshrimali5 :-) @Miracrown - You must have bought the ticket... :( It was really a nice movie.

ganeshie8 (ganeshie8):

I think she will join if the film is released in space theater : http://www.virgingalactic.com/ :)

OpenStudy (vishweshshrimali5):

We will have to look into that. Don't worry we will try to turn up with a solution for that also.

mathslover (mathslover):

A committee will be formed for this ... lol Now, as @Miracrown is not buying the ticket, unfortunately, the show ENDS here. lol

Miracrown (miracrown):

uhm- I will never buy the tickets-no matter what! even if it gets released in the space theater (which it never will, and its just virtual thinking... I will still not join. -.-